Skip to main content

AAVSOnet status, 2017-12-03

1 post / 0 new
HQA
HQA's picture
AAVSOnet status, 2017-12-03

Here is the current telescope status.

BSM_Berry:  fully operational

BSM_Hamren:  fully operational

BSM_NH:  fully operational.

BSM_NM:  the camera has moisture problems again.  Bill Stein is going to replace the Farpoint dessicant plug with the original SBIG plug and see if that helps.

BSM_South: fully operational.

SRO:  fully operational.

OC61:  fully operational.

TMO61:  the dome is still having communications problems, and the system is down until Jon Holtzman has the opportunity to work on the issue.

ongoing issues:

BSM_Hamren needs an automatic enclosure.  Right now, Mike Linnolt is using a tarp and the system can only be used for a few hours at the beginning of a night.

BSM_NH is having two minor issues.  Short exposures (say, 2 seconds or less) seem to intermittently hang during MaximDL download.  We haven't found the cause for this yet.  Scheduler is also moving to a target and waiting for a GEM flip before observing, and that flip can be an hour or two from when the field is centered, wasting a lot of time.  We need to understand the relationship between the software and the AVX mount hardware.

upcoming changes:

BSM_NH:  I will be adding a diffraction grating to the filter wheel.  This was a spare from BSM_Berry that was just returned to us by Greg Bolt.

BSM_South:  I will be sending a CFW-10 filter wheel plus new filters and a diffraction grating to Peter Nelson sometime soon.  Thanks to Frank Schorr and Tim Crawford for making this happen!

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484